5÷(√(x+1))-3√(x+1)+2=0


1
Expert's answer
2021-10-05T11:09:47-0400

Let "t=\sqrt{x+1}, t\geq0." Then


"\dfrac{5}{t}-3t+2=0"

"3t^2-2t-5=0"

"D=(-2)^2-4(3)(-5)=64"

"t=\dfrac{2\pm\sqrt{64}}{2(3)}=\dfrac{1\pm4}{3}"

Since "t\geq0," we take "t=\dfrac{1+4}{3}=\dfrac{5}{3}."

Then


"\sqrt{x+1}=\dfrac{5}{3}"

"x+1=\dfrac{25}{9}"

"x=\dfrac{16}{9}"